# WiFi-Diffusion: Achieving Fine-Grained WiFi Radio Map Estimation with Ultra-Low Sampling Rate by Diffusion Models

### v1

@@ Abstract
Fine-grained radio map presents communication parameters of interest, e.g., received signal strength, at every point across a large geographical region. It can be leveraged to improve the efficiency of spectrum utilization for a large area, particularly critical for the unlicensed WiFi spectrum. The problem of fine-grained radio map estimation is to utilize radio samples collected by sparsely distributed sensors to infer the map. This problem is challenging due to the ultra-low sampling rate, where the number of available samples is far less than the fine-grained resolution required for radio map estimation. We propose WiFi-Diffusion – a novel generative framework for achieving fine-grained WiFi radio map estimation using diffusion models. WiFi-Diffusion employs the creative power of generative AI to address the ultra-low sampling rate challenge and consists of three blocks: 1) a boost block, using prior information such as the layout of obstacles to optimize the diffusion model; 2) a generation block, leveraging the diffusion model to generate a candidate set of radio maps; and 3) an election block, utilizing the radio propagation model as a guide to find the best radio map from the candidate set. Extensive simulations demonstrate that 1) the fine-grained radio map generated by WiFi-Diffusion is ten times better than those produced by state-of-the-art (SOTA) when they use the same ultra-low sampling rate; and 2) WiFi-Diffusion achieves comparable fine-grained radio map quality with only one-fifth of the sampling rate required by SOTA.
@@ I Introduction
Radio maps depict parameters of interest in communication channels such as received signal strength (RSS) at every point of a certain geographical region [1]. They significantly enhance the performance of many wireless applications, enabling functions such as dynamic spectrum access [2], spectrum sharing [3], and interference management [4]. Fine-grained radio maps are radio maps of high resolutions for a large geographical region. They are particularly critical for WiFi, which operates in unlicensed spectra and plays a pivotal role in shaping the future of wireless connectivity. Currently, the WiFi spectra are extensively utilized by numerous devices, transmitting a vast amount of data. It has been estimated that almost $70\%$ of the global Internet traffic crosses a WiFi network [5]. With the anticipated rapid growth in WiFi traffic, advanced techniques in radio recognition, spectrum management, and network resource allocation are needed to ensure fast, reliable, and ubiquitous connectivity enabled by WiFi. Fine-grained radio maps can serve as an indispensable foundation for efficient utilization of WiFi in a large area.
Radio map estimation can generally be classified into two categories: sampling-free problems and sampling-based problems. Sampling-free problems use information of transmitters to estimate radio maps. Existing studies solving these problems include [6, 7, 8, 9]. Sampling-free problems assume the knowledge of information (e.g., number, locations, and transmitting power) of transmitters. This assumption is strong and may not hold in practice, especially for WiFi that operates in unlicensed spectra where there often exist many transmitters unknown to users. In contrast, sampling-based problems use certain prior information, e.g., layout of obstacles and limited number of radio samples collected by sensors sparsely distributed in the geographical region, to predict radio maps. In this paper, we focus on sampling-based fine-grained radio map estimation for the WiFi spectrum.
@@
Sampling-based radio map estimation approaches can generally be classified into two categories: interpolation-based approaches and deep learnning-based approaches. Interpolation-based approaches utilize mathematical and statistical methodologies to estimate radio maps. Well-known methods include radial basis functions (RBF) [10], spline [11], and ordinary kriging [12]. However, these methods often fail to construct high-quality radio maps for real-world environments that are dynamic and geographically complex. This is because they rely on mathematical calculations and do not account for prior information, such as layout of obstacles, which can significantly impact radio signal propagation.
@@
Recently, cutting-edge deep learning techniques have been applied to the problem of radio map estimation, surpassing the performance of traditional interpolation-based approaches. Levie et al. [13] proposed a U-Net-based scheme for estimating radio maps. Teganya et al. [14] and Locke et al. [15] designed autoencoder-based methods for constructing radio maps. He et al. [16] designed a ResNet-based approach for radio map estimation. Other deep learning approaches, e.g., Feedforward Neural Networks [17], Generative Adversarial networks [18], and Graph Attention Network [3], were used for radio map estimation, too.
@@
Fine-grained radio map estimation problem is challenging due to the ultra-low sampling rate, i.e., because of the inherent conflict between the fine-grained resolution required for radio map estimation and the ultra-small number of radio samples collected by sensors. In practice, the price of sensors that can collect radio samples is high. For instance, even a low-end sensor like Ettus USRP B200mini can cost over $1,000$ USD. To generate one radio map with a resolution of $256\times 256$ like those in [13] and [14], even a sampling rate of $1\%$ would require over $600$ sensors, resulting in a total cost exceeding $600,000$ USD, which is impractical. Therefore, it is crucial for radio map estimation algorithms to be capable of producing fine-grained, high-quality radio maps at ultra-low sampling rate (e.g., less than $0.1\%$). However, existing deep learning-based approaches, such as those in [15, 14, 16, 17, 18, 13, 3], have sampling rates ranging from $1\%$ to $10\%$. They all fail to construct high-quality radio maps at ultra-low sampling rates.
@@
In order to deal with the ultra-low sampling rate challenge, in this paper, we leverage advanced diffusion-based generative techniques for radio map estimation. One key advantage of generative models as compared to traditional discriminative models is its “out-of-the-box thinking” ability to generate a wide range of diverse predictions from limited inputs. This creativie power makes generative models well-suited for estimating radio maps at ultra-low sampling rates. Among the most prominent and powerful generative models is the diffusion model, which was first introduced in [19] and significantly improved in [20]. Since then, it has been successfully applied to various tasks[21] such as image synthesis [22], natural language processing [23], and molecule design [24]. To the best of our knowledge, we are the first to use diffusion-based generative techniques to solve the problem of radio map estimation. Specifically, we design WiFi-Diffusion, a diffusion model-based algorithm capable of generating fine-grained, high-quality radio maps at ultra-low sampling rates.
@@
As shown in Fig. 1, the key idea of WiFi-Diffusion is first to use the diffusion model to generate a candidate set of fine-grained radio maps and then find the best map from the set as the final output. To optimize the accuracy of the final output, WiFi-Diffusion solves the following two critical problems:
@@
To solve the first problem, WiFi-Diffusion incorporates a boost block, where it leverages prior information, including layout of obstacles, law of radio propagation, and collected radio samples, as inputs to a U-Net model with self-attention. The U-Net model with self-attention identifies “critical points” within the geographical region from the perspective of constructing radio maps and conveys the information of these points to the diffusion model. Subsequently, WiFi-Diffusion employs the diffusion model to generate a set of radio map candidates from the limited samples during a generation block. Although the qualities of candidate maps can be enhanced by the previous boost block, due to the ultra-low sampling rate, the majority of these maps are poor, with only a minority being accurate. Then WiFi-Diffusion proceeds to an election block that solves the second problem and selects the best map from the candidate set. In this block, WiFi-Diffusion uses the mathematical radio propagation model as a guide to identify the map that most closely adheres to the propagation model in the candidate set and selects it as the final output. Through this novel generative framework of three blocks, WiFi-Diffusion is able to generate fine-grained, high-quality radio maps at ultra-low sampling rates. The main contributions of this paper are summarized as follows:
@@
  • We propose WiFi-Diffusion – a novel generative framework for fine-grained radio map estimation for the WiFi spectrum. To address the ultra-low sampling rate challenge of fine-grained radio map estimation, WiFi-Diffusion leverages the creative power of diffusion-based generative model to produce a diverse set of radio map candidates. Subsequently, WiFi-Diffusion identifies and selects the best map from this set as the final output. To the best of our knowledge, we are the first to use diffusion-based generative techniques to solve the problem of sampling-based radio map estimation.
@@
  • Although a diverse set of radio maps can be generated by the diffusion model, the majority of them are quite poor, with only a minority being accurate. To guarantee a high-quality final output, WiFi-Diffusion incorporates a boost block prior to generating radio map candidates and an election block following their generation. In the boost block, WiFi-Diffusion utilizes prior information, e.g., layout of obstacles, principle of radio propagation, and collected radio samples, to enhance the qualities of radio map candidates generated by the diffusion model. In the election block, WiFi-Diffusion employs the mathematical radio propagation model as a guide to find the best map from the diverse candidates generated by the diffusion model. With the boost block, generation block, and election block, WiFi-Diffusion is able to generate high-quality, fine-grained radio maps at ultra-low sampling rates.
@@
  • Extensive simulations based on the 5750 MHz (WiFi 5G) BART-Lab radio map dataset show that WiFi-Diffusion is effective. Specifically, simulations demonstrate that 1) WiFi-Diffusion generates high-quality, fine-grained radio maps when the sampling rate is less than $0.1\%$; 2) the radio map generated by WiFi-Diffusion is ten times better than that generated by state-of-the-art (SOTA) when they share the same ultra-low sampling rate; 3) WiFi-Diffusion requires only one-fifth of the sampling rate needed by SOTA to generate comparable radio maps.
@@
The rest of the paper is organized as follows. In Section II, we describe the system model and introduce the radio map estimation problem. In Section III, we present the key ideas of the design of WiFi-Diffusion. In Section IV, we offer the design details of WiFi-Diffusion. In Section V, we perform extensive simulations to evaluate the performance of WiFi-Diffusion. Section VI concludes the paper.
@@ II Problem Description
Consider a two-dimensional region of interest $\mathcal{Y}\subset\mathbb{R}^{2}$. Consider a set of sensors $\mathcal{S}=\{1,2,\cdots,k\}$ (the number of sensors is $k$) where each sensor $s\in\mathcal{S}$ is located at a unique coordinate $y_{s}\in\mathcal{Y}$. Each sensor $s$ measures the RSS in a given WiFi frequency band $\mathcal{F}$ at its location $y_{s}$, where $r_{s}(f)$ denotes the RSS in frequency $f\in\mathcal{F}$ (assuming that effects of fast-fading are averaged out across minor temporal or spatial shifts during the measurement). These radio samples $r_{s}(f)$ ($s\in\mathcal{S}$ and $f\in\mathcal{F}$) are reported to a fusion center, which may be e.g. a base station or a cloud server, depending on the application.
To facilitate mathematical analysis, in this paper, we narrow our focus to a single frequency $f$, considering only spatial samples with RSS $r_{s}$ for each $s\in\{1,2,\cdots,k\}$ at corresponding location $y_{s}\in\mathcal{Y}$. Moreover, we discretize the continuous space $\mathcal{Y}$ into a $H\times W$ grid (with height $H$ and width $W$). Hence, each location $y_{s}$ corresponds to a grid index $(h_{s},w_{s})$ where $h_{s}\in\{0,1,\cdots,H-1\}$ and $w_{s}\in\{0,1,\cdots,W-1\}$. Define $\mathbf{S}_{k}$ as an $H\times W$ matrix, where the element at index $(h_{s},w_{s})$ is 0 if there is no sensor located at that grid point; otherwise, it is the RSS value $r_{s}$ of the sample collected by sensor $s$ with location $(h_{s},w_{s})$.
To ground our findings in reality, we take environmental obstacles, specifically buildings, into consideration. These structures can occupy any grid point within the region $\mathcal{Y}$, often covering multiple adjacent points. We consider the occupied grid points to have no RSS, however they do affect RSS in surrounding areas by reflecting radio waves back into the environment and by blocking or significantly attenuating radio waves passing through them. The layout of buildings in this paper is represented by a matrix $\mathbf{B}$, where $\mathbf{B}(h_{s},w_{s})\in\{0,1\}$ with $\mathbf{B}(h_{s},w_{s})=0$ denoting that there is no building at the grid point $(h_{s},w_{s})$ and $\mathbf{B}(h_{s},w_{s})=1$ otherwise.
@@
The problem of radio map estimation is for the fusion center is to use $\mathbf{S}_{k}$ and $\mathbf{B}$ to infer (estimate) the RSS at every grid point within the region $\mathcal{Y}$. As shown in Fig. 2, the problem of radio map estimation requires the fusion center to use the information of Fig. 2a ($\mathbf{S}_{k}$, which gives RSS values of collected radio samples) and Fig. 2b ($\mathbf{B}$, which is the layout of buildings) to infer Fig. 2c (RSS values at each location in the region of interest).
@@ III WiFi-Diffusion: Basic Ideas
The problem of fine-grained radio map estimation is extremely challenging due to that the number of samples gathered by sensors is extremely limited in practice. It is almost impossible for traditional algorithms to construct a radio map when only a tiny fraction, e.g., $0.1\%$, of the points in the map are known. To overcome this challenge, we propose WiFi-Diffusion -— an algorithm that leverages diffusion-based generative techniques to construct fine-grained, high-quality radio maps at ultra-low sampling rates.
As shown in Fig. 3, WiFi-Diffusion consists of three blocks: a boost block, a generation block, and an election block. In the following, we briefly describe these three blocks. The objective of the boost block is to optimize and enhance the subsequent generation block. It would be impossible to discuss the boost block without gaining a clear understanding of the generation block first. So we first describe the generation block.
Block 2: Generation block    In this block, WiFi-Diffusion uses the diffusion-based generative model Denoising Diffusion Implicit Model (DDIM) to produce a candidate set of radio maps from the limited known radio samples. Generative models have an “out-of-the-box thinking” ability and can make diverse predictions (despite that majority of which may not be accurate) from extremely limited inputs. For the problem of radio map estimation, we use DDIM to generate multiple (64 in simulations) radio maps from only a few (less than $0.1\%$ in simulations) radio samples. While the majority of these generated maps may deviate significantly from the true radio environment, we expect that there are a few high-quality maps within the set.
@@
Block 1: Boost block    Although generative models like DDIM exhibit creativity, the extremely limited number of radio samples often results in a candidate set of maps that are predominantly or entirely of low quality. Therefore, it is important to design a boost block preceding the generation block, improving the quality of the maps in the candidate set. Specifically, in the boost block, WiFi-Diffusion leverages prior information, including layout of buildings, principle of radio propagation, and collected radio samples, as inputs to a U-Net model with self-attention. The U-Net model with self-attention identifies “critical points” within the geographical region $\mathcal{Y}$ from the perspective of constructing radio maps and conveys the information of these points to DDIM for use in the subsequent generation block.
Block 3: Election block    Even with the boost block, many maps in the candidate set generated by DDIM are of low-quality. Therefore, it is necessary to design an election block following the generation block, finding the best map in the candidate set. Specifically, in the election block, WiFi-Diffusion employs the mathematical radio propagation model as a guide to identify the map that most closely adheres to the propagation model in the candidate set. WiFi-Diffusion selects it as the best map and considers it as the final output.
With the above-mentioned three blocks, WiFi-Diffusion is able to construct fine-grained, high-quality radio maps when the sampling rate is extremely low. In the following section, we describe details of the three blocks of WiFi-Diffusion.

### v3

@@ Abstract
The radio map presents communication parameters of interest, e.g., received signal strength, at every point across a geographical region. It can be leveraged to improve the efficiency of spectrum utilization in the region, particularly critical for unlicensed WiFi spectrum. The problem of fine-grained radio map estimation is to utilize radio samples collected by sensors sparsely distributed in the region to infer a high-resolution radio map. This problem is challenging due to the ultra-low sampling rate, i.e., because the number of available samples is far fewer than the high resolution required for radio map estimation. We propose WiFi-Diffusion – a novel generative framework for achieving fine-grained WiFi radio map estimation using diffusion models. WiFi-Diffusion employs the creative power of generative AI to address the ultra-low sampling rate challenge and consists of three blocks: 1) a boost block, using prior information such as the layout of obstacles to optimize the diffusion model; 2) a generation block, leveraging the diffusion model to generate a candidate set of fine-grained radio maps; and 3) an election block, utilizing the radio propagation model as a guide to find the best fine-grained radio map from the candidate set. Extensive simulations demonstrate that 1) the fine-grained radio map generated by WiFi-Diffusion is ten times better than those produced by state-of-the-art (SOTA) when they use the same ultra-low sampling rate; and 2) WiFi-Diffusion achieves comparable fine-grained radio map quality with only one-fifth of the sampling rate required by SOTA.
@@ I Introduction
The radio map depicts parameters of interest in communication channels, such as the received signal strength (RSS), at every point of a certain geographical region [1]. It can significantly improve the performance of many wireless applications, enabling functions such as dynamic spectrum access [2], spectrum sharing [3], and interference management [4]. The fine-grained radio map is a high-resolution radio map for a large geographical region. It is particularly critical for WiFi, which operates in unlicensed spectra and plays a pivotal role in shaping the future of wireless connectivity. Currently, the WiFi spectra are extensively utilized by numerous devices, transmitting a vast amount of data. It has been estimated that almost $70\%$ of global Internet traffic crosses a WiFi network [5]. With the anticipated rapid growth in WiFi traffic, advanced techniques in radio recognition, spectrum management, and network resource allocation are needed to ensure fast, reliable, and ubiquitous connectivity enabled by WiFi. The fine-grained radio map can serve as an indispensable foundation for the efficient utilization of WiFi in a large area.
Radio map estimation problems can generally be classified into two categories: sampling-free problem and sampling-based problem. The sampling-free problem uses the information from the transmitters to estimate radio maps. Existing studies solving these problems include [6, 7, 8, 9]. The sampling-free problem assumes knowledge of the information (e.g., number, locations, and transmitting power) of transmitters. This assumption is strong and may not hold in practice, especially for WiFi that operates in unlicensed spectra, where there often exist many transmitters unknown to users. In contrast, the sampling-based problem uses certain prior information, e.g., the layout of obstacles and the limited number of radio samples collected by sensors sparsely distributed in the geographical region, to predict radio maps. In this paper, we focus on the sampling-based fine-grained radio map estimation problem for the WiFi spectrum.
@@
Approaches for solving the sampling-based radio map estimation problem can generally be classified into two categories: interpolation-based approaches and deep learning-based approaches. Interpolation-based approaches utilize mathematical and statistical methodologies to estimate radio maps. Well-known methods include radial basis functions (RBF) [10], spline [11], and ordinary kriging [12]. However, these methods often fail to construct high-quality radio maps for real-world environments that are dynamic and geographically complex. This is because they rely on mathematical calculations and do not account for certain critical prior information, such as the layout of obstacles, which can significantly impact radio signal propagation.
@@
Recently, cutting-edge deep learning techniques have been applied to the problem of radio map estimation, surpassing the performance of traditional interpolation-based approaches. Levie et al. [13] proposed a U-Net-based scheme for estimating radio maps. Teganya et al. [14] and Locke et al. [15] designed autoencoder-based methods for constructing radio maps. He et al. [16] designed a ResNet-based approach for radio map estimation. Other deep learning approaches, e.g., Feedforward Neural Networks [17], Generative Adversarial networks [18], and Graph Attention Networks [3], were used for radio map estimation, too.
@@
The fine-grained radio map estimation problem is challenging due to the ultra-low sampling rate, i.e., because of the inherent conflict between the high resolution required for radio map estimation and the ultra-small number of radio samples collected by sensors. In practice, the price of sensors that can collect radio samples is high. For example, even a low-end sensor like the Ettus USRP B200mini can cost more than $1,000$ USD. To generate a radio map with a resolution of $256\times 256$ like those in [13] and [14], even a sampling rate of $1\%$ would require over $600$ sensors, resulting in a total cost exceeding $600,000$ USD, which is impractical. Therefore, it is crucial for radio map estimation approaches to be capable of producing fine-grained, high-quality radio maps at ultra-low sampling rates (e.g., less than $0.1\%$). However, existing deep learning-based approaches, such as those in [15, 14, 16, 17, 18, 13, 3], have sampling rates ranging from $1\%$ to $10\%$. They all fail to construct fine-grained, high-quality radio maps at ultra-low sampling rates.
@@
In order to deal with the ultra-low sampling rate challenge, in this paper, we leverage advanced diffusion-based generative techniques for radio map estimation. One key advantage of generative models compared to traditional discriminative models is their “out-of-the-box thinking” ability to generate a wide range of diverse predictions from limited inputs. This creative power makes generative models well-suited for estimating radio maps at ultra-low sampling rates. Among the most prominent and powerful generative models is the diffusion model, which was first introduced in [19] and significantly improved in [20]. Since then, it has been successfully applied to various tasks such as image synthesis [21], natural language processing [22], and molecule design [23]. To the best of our knowledge, we are the first to use diffusion-based generative techniques to solve the problem of radio map estimation. Specifically, we design WiFi-Diffusion, a diffusion model-based algorithm capable of generating fine-grained, high-quality radio maps at ultra-low sampling rates.
@@
As shown in Fig. 1, the key idea of WiFi-Diffusion is first to use the diffusion model to generate a candidate set of fine-grained radio maps and then to find the best map from the set as the final output. To optimize the quality of the final output, WiFi-Diffusion solves the following two critical problems:
@@
To solve the first problem, WiFi-Diffusion incorporates a boost block, where it leverages prior information, including the layout of obstacles, the law of radio propagation, and collected radio samples, as inputs to a UNet model with self-attention. The UNet model identifies “critical points” within the geographical region from the perspective of constructing radio maps and conveys the information of these points to the diffusion model. Subsequently, WiFi-Diffusion employs the diffusion model to generate a set of fine-grained radio map candidates from the limited samples during a generation block. Although the qualities of candidate maps can be enhanced by the previous boost block, due to the ultra-low sampling rate, the majority of these maps are poor, with only a minority being accurate. Then WiFi-Diffusion proceeds to an election block that solves the second problem and selects the best map from the candidate set. In this block, WiFi-Diffusion uses the mathematical radio propagation model as a guide to identify the map that most closely adheres to the propagation model in the candidate set and selects it as the final output. Through this novel generative framework of three blocks, WiFi-Diffusion is able to generate fine-grained, high-quality radio maps at ultra-low sampling rates.
Specifically, the main contributions of this paper are summarized as follows:
@@
  • We propose WiFi-Diffusion – a novel generative framework for fine-grained radio map estimation for the WiFi spectrum. To address the ultra-low sampling rate challenge of fine-grained radio map estimation, WiFi-Diffusion leverages the creative power of diffusion-based generative models to produce a diverse set of fine-grained radio map candidates. Subsequently, WiFi-Diffusion identifies and selects the best map from this set as the final output. To the best of our knowledge, we are the first to use diffusion-based generative techniques to solve the problem of sampling-based radio map estimation.
@@
  • Although a diverse set of fine-grained radio maps can be generated by the diffusion model, the majority of them are quite poor and only a minority are accurate. To guarantee a high-quality final output, WiFi-Diffusion incorporates a boost block prior to generating fine-grained radio map candidates and an election block following their generation. In the boost block, WiFi-Diffusion utilizes prior information, e.g., layout of obstacles, principle of radio propagation, and collected radio samples, to enhance the qualities of fine-grained radio map candidates generated by the diffusion model. In the election block, WiFi-Diffusion employs the mathematical radio propagation model as a guide to find the best map from the diverse candidates generated by the diffusion model. With the boost block, generation block, and election block, WiFi-Diffusion is able to generate fine-grained, high-quality radio maps at ultra-low sampling rates.
@@
  • Extensive simulations based on the 5750 MHz (WiFi 5G) BART-Lab radio map dataset show that WiFi-Diffusion is effective. Specifically, simulations demonstrate that 1) WiFi-Diffusion generates fine-grained, high-quality radio maps when the sampling rate is less than $0.1\%$; 2) the fine-grained radio map generated by WiFi-Diffusion is ten times better than that generated by state-of-the-art (SOTA) when they share the same ultra-low sampling rate; 3) WiFi-Diffusion requires only one-fifth of the sampling rate needed by SOTA to generate comparable fine-grained radio maps.
@@
The rest of the paper is organized as follows. In Section II, we describe the system model and introduce the radio map estimation problem. In Section III, we present the key ideas of the design of WiFi-Diffusion. In Section IV, we offer the design details of WiFi-Diffusion. In Section V, we perform extensive simulations to evaluate the performance of WiFi-Diffusion. Section VI discusses potential applications of WiFi-Diffusion. Section VII concludes the paper.
@@ II Problem Description
Consider a two-dimensional region of interest $\mathcal{Y}\subset\mathbb{R}^{2}$. Consider a set of sensors $\mathcal{S}=\{1,2,\cdots,k\}$ (the number of sensors is $k$) where each sensor $s\in\mathcal{S}$ is located at a unique coordinate $y_{s}\in\mathcal{Y}$. Each sensor $s$ measures the RSS in a given WiFi frequency band $\mathcal{F}$ at its location $y_{s}$, where $r_{s}(f)$ denotes the RSS in frequency $f\in\mathcal{F}$ (assuming that the effects of fast-fading are averaged out across minor temporal or spatial shifts during the measurement). These radio samples $r_{s}(f)$ ($s\in\mathcal{S}$ and $f\in\mathcal{F}$) are reported to a fusion center, which may be, e.g., a base station or a cloud server, depending on the application.
To facilitate mathematical analysis, in this paper, we narrow our focus to a single frequency $f$, considering only spatial samples with RSS $r_{s}$ for each $s\in\{1,2,\cdots,k\}$ at the corresponding location $y_{s}\in\mathcal{Y}$. Moreover, we discretize the continuous space $\mathcal{Y}$ into a $H\times W$ grid (with height $H$ and width $W$). Hence, each location $y_{s}$ corresponds to a grid index $(h_{s},w_{s})$ where $h_{s}\in\{0,1,\cdots,H-1\}$ and $w_{s}\in\{0,1,\cdots,W-1\}$. Define $\mathbf{S}_{k}$ as an $H\times W$ matrix, where its element at index $(h_{s},w_{s})$, i.e., $\mathbf{S}_{k}(h_{s},w_{s})$, is 0 if there is no sensor located at that grid point; otherwise, it is the RSS value $r_{s}$ of the sample collected by the sensor $s$ with location $(h_{s},w_{s})$.
To ground our findings in reality, we take environmental obstacles, specifically buildings, into consideration. These structures can occupy any grid point within the region $\mathcal{Y}$, often covering multiple adjacent points. We consider the occupied grid points to have no RSS; however, they do affect RSS in surrounding areas by reflecting radio waves back into the environment and by blocking or significantly attenuating radio waves passing through them. The layout of buildings in this paper is represented by a matrix $\mathbf{B}$, where its element at index $(h_{s},w_{s})$, i.e., $\mathbf{B}(h_{s},w_{s})$, is 00 if there is no building at the point of the grid $(h_{s},w_{s})$; otherwise, it is $1$.
@@
The problem of radio map estimation requires the fusion center to use $\mathbf{S}_{k}$ and $\mathbf{B}$ to infer (estimate) the RSS at every grid point within the region $\mathcal{Y}$. Specifically, as shown in Fig. 2, the problem of radio map estimation requires the fusion center to use the information of Fig. 2a ($\mathbf{S}_{k}$, which gives RSS values of collected radio samples) and Fig. 2b ($\mathbf{B}$, which is the layout of buildings) to infer Fig. 2c (RSS values at each location in the region of interest).
@@ III WiFi-Diffusion: Basic Ideas
The problem of fine-grained radio map estimation is extremely challenging because the number of samples collected by the sensors is extremely small compared to the high resolution of radio maps to be estimated in practice. It is almost impossible for traditional approaches to construct a radio map when only a tiny fraction, e.g., $0.1\%$, of the points in the map are known. To overcome this challenge, we propose WiFi-Diffusion -— an algorithm that leverages diffusion-based generative techniques to construct fine-grained, high-quality radio maps at ultra-low sampling rates.
As shown in Fig. 3, WiFi-Diffusion consists of three blocks: a boost block, a generation block, and an election block. In the following, we briefly describe these three blocks. The objective of the boost block is to optimize and enhance the subsequent generation block. It would be impossible to discuss the boost block without first gaining a clear understanding of the generation block. So we first describe the generation block.
Block 2: Generation block    In this block, WiFi-Diffusion uses the diffusion-based generative model Denoising Diffusion Implicit Model (DDIM) to produce a candidate set of fine-grained radio maps from the limited known radio samples. Generative models have an “out-of-the-box thinking” ability and can make diverse predictions (despite that the majority of which may not be accurate) from extremely limited inputs. For the problem of fine-grained radio map estimation, we use DDIM to generate multiple fine-grained radio maps (64 in simulations) from only a few (less than $0.1\%$ in simulations) radio samples. Although most of these generated maps may deviate significantly from the true radio environment, we expect that there are a few high-quality maps within the set.
@@
Block 1: Boost block    Although generative models like DDIM exhibit creativity, the extremely limited number of radio samples often results in a candidate set of maps that are predominantly or entirely of low quality. Therefore, it is important to design a boost block preceding the generation block, improving the quality of the maps in the candidate set. Specifically, in the boost block, WiFi-Diffusion leverages prior information, including the layout of buildings, principles of radio propagation, and collected radio samples, as inputs to a UNet model with self-attention. The UNet model identifies “critical points” within the geographical region $\mathcal{Y}$ from the perspective of constructing radio maps and conveys the information of these points to DDIM for use in the subsequent generation block.
Block 3: Election block    Even with the boost block, many maps in the candidate set generated by DDIM are of low quality. Therefore, it is necessary to design an election block following the generation block to find the best map in the candidate set. Specifically, in the election block, WiFi-Diffusion employs the mathematical radio propagation model as a guide to identify the map that most closely adheres to the propagation model in the candidate set. WiFi-Diffusion selects it as the best map and considers it as the final output.
With the above-mentioned three blocks, WiFi-Diffusion is able to construct fine-grained, high-quality radio maps when the sampling rate is extremely low. In the following section, we describe the details of the three blocks of WiFi-Diffusion.